Let : 

V = R3, 

W = {(x1, X2, X3) ! x1 - x2 = .X3}. 

Show that W is a subspace of V. Further, 

find a basis for W, and hence, find the 

dimension of W.


1
Expert's answer
2021-02-12T10:32:28-0500

Let "V = \mathbb R^3""W = \{(x_1, x_2, x_3) |\ x_1 - x_2 = x_3\}"

Let us show that "W" is a subspace of "V". Let "(x_1,x_2,x_3), (y_1,y_2,y_3)\in W," and "\ a, b\in\mathbb R". Then "x_1-x_2=x_3" and "y_1-y_2=y_3". It follows that "(ax_1+by_1)-(ax_2+by_2)=a(x_1-x_2)+b(y_1-y_2)=ax_3+by_3" , and therefore, "a(x_1,x_2,x_3)+ b(y_1,y_2,y_3)= (ax_1+by_1,ax_2+by_2,ax_3+by_3)\in W". We conclude that "W" is a subspace of "V".


Further, find a basis for "W" . Since "W = \{(x_1, x_2, x_3)\in\mathbb R^3 |\ x_1 - x_2 = x_3\}= \{(x_1, x_2, x_1-x_2)\in\mathbb R^3 |\ x_1, x_2\in\mathbb R\}= \{x_1(1, 0, 1)+x_2(0, 1, -1)\in\mathbb R^3 |\ x_1, x_2\in\mathbb R\},"we conclude that linear independent vectors "e_1=(1, 0, 1)" and "e_2=(0, 1, -1)" form the basis of "W",

and hence "\dim(W)=2."